OpenStudy (anonymous):

how do I evaluate: lim x->0 (sin^2(x))/(1-cosx)??

OpenStudy (anonymous):

use lehopitals rule difrentiat top and botom twice then plug in limit

OpenStudy (anonymous):

LOL omg i cant believe

OpenStudy (anonymous):

no need for that

OpenStudy (anonymous):

i deleted it by accident :( but : lhosiptal as i said 2sinxcosx / sinx 2cosx plug x = 0 2

OpenStudy (anonymous):

\[\frac{(1-\cos(x))(1+\cos(x))}{1-\cos(x)}=1+\cos(x)\]

OpenStudy (anonymous):

haha i got it before you deleted it. I kinda understand what you are all saying? but im still a little confused and have more evaluations to do...

OpenStudy (anonymous):

replace x by 0, get 2 right away

myininaya (myininaya):

nice eye satellite

OpenStudy (anonymous):

look at the way i solved it now .. you dont really need to do it twice as i did before

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

anyone know a good site for trigonometric identities? thats what gets me

OpenStudy (anonymous):

paul's notes is alway good but if you need one on the fly just google it

OpenStudy (anonymous):

could you possibly help me with another? i think i may have gotten the first step but i am not sure

OpenStudy (anonymous):

sure post away